1. How much energy is required to melt a 0.5kg block of ice?





  1. Calculate the specific latent heat of fusion of naphthalene given that 6x105J of heat are given out when 4.0kg of naphthalene at its melting point changes to a solid.





  1. How much energy and time does a 2200w kettle require to evaporate 0.5kg of water at 1000C? (Specific latent heat of vaporization of water = 2.26x106J/kg)





  1. Calculate the amount of heat energy required to melt 0.3 kg of ice at 0 °C.

(Specific latent heat of fusion of ice = 3.34 x 10J/kg)



  1. What is the difference between latent heat of fusion and latent heat of vaporisation of a substance?
1
Expert's answer
2021-05-26T10:33:17-0400

1.

"Q=\lambda m=3.34\cdot 10^5\cdot 0.5=167~kJ."

2.

"\lambda=\frac Qm=\frac{6\cdot 10^5}{4}=1.5\cdot 10^5~\frac {J}{kg}."

3.

"P=\frac Qt\implies t=\frac QP=\frac{Lm}{P}=\frac{2.26\cdot 10^6\cdot 0.5}{2200}=8.5~min."

4.

"Q=\lambda m=3.34\cdot 10^5\cdot 0.3=0.1~MJ."

5.

The difference in the transition to the liquid and gaseous phase, respectively.